# Distributional Limits for Eigenvalues of Graphon Kernel Matrices

### v1

@@ 6 Proof of the main theorem
This section contains the proof of Theorem 3.3. We follow the strategy outlined in Section 3.1: first reduce the eigenvalue fluctuation to a Rayleigh quotient (up to a negligible error) using perturbation theory for simple eigenvalues, and then analyze the resulting quadratic form via a Hoeffding decomposition. The argument splits into two regimes depending on whether the linear Hoeffding projection is nontrivial ($\sigma_{r}^{2}>0$) or vanishes identically ($\sigma_{r}^{2}=0$).
Proof of Theorem 3.3.
@@
By Lemma A.0.3 and the eigengap condition $\gamma_{r}>0$ for $T_{W}$, the eigenvalues of $\mathbf{K}_{n}/(n-1)$ can be indexed so that $\lambda_{k}(\mathbf{K}_{n})/(n-1)\xrightarrow{p}\lambda_{k}$ for each fixed $k$. In particular, with probability tending to $1$, all eigenvalues of $\mathbf{K}_{n}$ other than $\lambda_{r}(\mathbf{K}_{n})$ lie at distance at least $cn$ from $(n-1)\lambda_{r}$ for some $c>0$. On this event, set
@@
  • $\theta:=\mathbb{E}[h_{r}(U,U^{\prime})]=\lambda_{r}$ (as computed in the proof of Lemma B.2.6);
@@
Let $\mathbf{V}\in\mathbb{R}^{n\times(n-1)}$ have orthonormal columns spanning $\mathbf{u}^{\perp}$. By standard perturbation theory for simple eigenvalues (or Lemma A.0.5), the eigenvalue $\lambda_{r}(\mathbf{K}_{n})$ then admits the expansion
@@
and the Hoeffding decomposition $h_{r}=h_{r,1}+h_{r,2}$ with $h_{r,1}\equiv\lambda_{r}$ (Proof of Lemma B.2.6) yields
@@
For the second term in (17), Lemma A.0.4 gives
@@
Using the joint convergence from the proof of Lemma B.2.6, for any fixed $K$ the vector $(T_{k,n})_{k\leq K,k\neq r}$ converges jointly in distribution to $(Z_{k})_{k\leq K,k\neq r}$, where $(Z_{k})$ are i.i.d. $\mathcal{N}(0,1)$. By the Hilbert-Schmidt condition $\sum_{k\geq 1}\lambda_{k}^{2}<\infty$ and the eigengap bound $|\lambda_{r}-\lambda_{k}|\geq\gamma_{r}$ for all $k\neq r$,
@@
converges in $L^{2}$. Together with the $L^{2}$ tail control in Lemmas B.2.6 and A.0.4, this implies
@@
Throughout this appendix, Assumption 3.1 is in force. Recall that $W\in L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$, that $T_{W}$ is the associated integral operator, and that $(\lambda_{k},\varphi_{k})$ denote its eigenpairs.
Lemma A.0.1 (Approximate eigenpair for the target index $r$).
Under Assumption 3.1, define
@@
Lemma A.0.2 (Kato-Temple inequality for symmetric matrices).
@@
This is a finite-dimensional specialization of the Kato-Temple inequality for self-adjoint operators; see Kato Kato (1995, Theorems V.4.10-V.4.11). ∎
Lemma A.0.3 (Kernel matrix eigenvalue scaling in (n1)(n-1) normalization).
Under Assumption 3.1, let
@@
Let $T_{W}$ be the integral operator on $L^{2}([0,1])$ with kernel $W$ and eigenvalues $\lambda_{1}(T_{W})\geq\lambda_{2}(T_{W})\geq\cdots$. Fix $r$ as in Assumption 3.1. Then
@@
Lemma A.0.4 (Resolvent correction in the degenerate regime).
@@
Fix the approximate eigenvectors $\mathbf{u}_{k}$ as in Lemma A.0.1, and for each fixed $K$ consider the family $\{\mathbf{u}_{k}\}_{k\leq K,k\neq r}$, which is asymptotically orthonormal and lies in $\mathbf{u}^{\perp}$ up to $o_{p}(1)$ by Lemma A.0.1. By Gram-Schmidt, choose $\mathbf{V}$ so that, for each fixed $K$, its first K1K-1 columns agree with an orthonormalization of $\{\mathbf{u}_{k}\}_{k\leq K,k\neq r}$, noting that the quadratic form
@@
depends only on the subspace $\mathbf{u}^{\perp}$ and $A$, not on the specific orthonormal basis $\mathbf{V}$, so this choice is without loss of generality. By Lemma A.0.3 and the simple eigengap at $\lambda_{r}$, the eigenvalues of $(n-1)^{-1}\mathbf{K}_{n}$ in $\mathbf{u}^{\perp}$ concentrate around $\{\lambda_{k}\}_{k\neq r}$, and hence
@@
so that $s_{k,n}=n(1+o_{p}(1))$ by Lemma B.2.7, and Lemma A.0.1 gives $\mathbf{u}_{k}^{\top}\mathbf{u}_{\ell}=\delta_{k\ell}+o_{p}(1)$ and
@@
where we used lemma  A.0.1 for $\mathbf{U}^{\top}\mathbf{U}$ and $\mathbf{M}_{K}$. For $k\leq K$, $k\neq r$, the approximate eigenvalue equation yields
@@
Lemma A.0.5 (Second-order Rayleigh-Schrödinger expansion for a simple eigenvalue).
@@
Lemma B.2.6 (Degenerate U-statistic in the degenerate regime).
@@
Lemma B.2.7.

### v2

@@ 6 Proof of the main theorem
This section contains the proof of Theorem 3. We follow the strategy outlined in Section 3.1: first reduce the eigenvalue fluctuation to a Rayleigh quotient (up to a negligible error) using perturbation theory for simple eigenvalues, and then analyze the resulting quadratic form via a Hoeffding decomposition. The argument splits into two regimes depending on whether the linear Hoeffding projection is nontrivial ($\sigma_{r}^{2}>0$) or vanishes identically ($\sigma_{r}^{2}=0$).
Proof of Theorem 3.
@@
By Lemma 10 and the eigengap condition $\gamma_{r}>0$ for $T_{W}$, the eigenvalues of $\mathbf{K}_{n}/(n-1)$ can be indexed so that $\lambda_{k}(\mathbf{K}_{n})/(n-1)\xrightarrow{p}\lambda_{k}$ for each fixed $k$. In particular, with probability tending to $1$, all eigenvalues of $\mathbf{K}_{n}$ other than $\lambda_{r}(\mathbf{K}_{n})$ lie at distance at least $cn$ from $(n-1)\lambda_{r}$ for some $c>0$. On this event, set
@@
  • $\theta:=\mathbb{E}[h_{r}(U,U^{\prime})]=\lambda_{r}$ (as computed in the proof of Lemma 13);
@@
Let $\mathbf{V}\in\mathbb{R}^{n\times(n-1)}$ have orthonormal columns spanning $\mathbf{u}^{\perp}$. By standard perturbation theory for simple eigenvalues (or Lemma 12), the eigenvalue $\lambda_{r}(\mathbf{K}_{n})$ then admits the expansion
@@
and the Hoeffding decomposition $h_{r}=h_{r,1}+h_{r,2}$ with $h_{r,1}\equiv\lambda_{r}$ (Proof of Lemma 13) yields
@@
For the second term in (17), Lemma 11 gives
@@
Using the joint convergence from the proof of Lemma 13, for any fixed $K$ the vector $(T_{k,n})_{k\leq K,k\neq r}$ converges jointly in distribution to $(Z_{k})_{k\leq K,k\neq r}$, where $(Z_{k})$ are i.i.d. $\mathcal{N}(0,1)$. By the Hilbert-Schmidt condition $\sum_{k\geq 1}\lambda_{k}^{2}<\infty$ and the eigengap bound $|\lambda_{r}-\lambda_{k}|\geq\gamma_{r}$ for all $k\neq r$,
@@
converges in $L^{2}$. Together with the $L^{2}$ tail control in Lemmas 13 and 11, this implies
@@
Acknowledgments
The author thanks Mohammad S. M. Moakhar for insightful discussions and valuable suggestions.
@@ Acknowledgments
Throughout this appendix, Assumption 1 is in force. Recall that $W\in L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$, that $T_{W}$ is the associated integral operator, and that $(\lambda_{k},\varphi_{k})$ denote its eigenpairs.
Lemma 7 (Approximate eigenpair for the target index $r$).
Under Assumption 1, define
@@
Lemma 8 (Sample eigengap bound).
Under Assumption 1, let $T_{W}$ denote the integral operator with eigenvalues $(\lambda_{k})_{k\geq 1}$ satisfying the eigengap $\gamma_{r}=\inf_{k\neq r}|\lambda_{r}-\lambda_{k}|>0$. Let $\mathbf{K}_{n}$ be the kernel matrix defined in Lemma 10, and denote its ordered eigenvalues by $\lambda_{1}(\mathbf{K}_{n})\geq\lambda_{2}(\mathbf{K}_{n})\geq\cdots\geq\lambda_{n}(\mathbf{K}_{n})$. Then there exists a constant $c>0$ such that
$\mathbb{P}\!\left(\min_{k\neq r}\big|\lambda_{r}(\mathbf{K}_{n})-\lambda_{k}(\mathbf{K}_{n})\big|\geq cn\right)\to 1.$
Proof.
From Lemma 10,
$\Big\|\frac{1}{n-1}\mathbf{K}_{n}-T_{W}\Big\|_{\mathrm{op}}=O_{p}(n^{-1/2}).$
By Weyl’s inequality for self-adjoint matrices,
$\Big|\frac{\lambda_{k}(\mathbf{K}_{n})}{n-1}-\lambda_{k}(T_{W})\Big|\leq\Big\|\frac{1}{n-1}\mathbf{K}_{n}-T_{W}\Big\|_{\mathrm{op}}.$
Hence, with probability 1o(1)1-o(1),
$\min_{k\neq r}\Big|\frac{\lambda_{r}(\mathbf{K}_{n})}{n-1}-\frac{\lambda_{k}(\mathbf{K}_{n})}{n-1}\Big|\geq\frac{\gamma_{r}}{2}.$
Multiplying both sides by (n1)(n-1) gives
$\min_{k\neq r}|\lambda_{r}(\mathbf{K}_{n})-\lambda_{k}(\mathbf{K}_{n})|\geq cn$
for some $c=\gamma_{r}/4$, completing the proof. ∎
Lemma 9 (Kato-Temple inequality for symmetric matrices).
@@
This is a finite-dimensional specialization of the Kato-Temple inequality for self-adjoint operators; see  Kato (1995, Theorems V.4.10-V.4.11). ∎
Lemma 10 (Kernel matrix eigenvalue scaling in (n1)(n-1) normalization).
Under Assumption 1, let
@@
Let $T_{W}$ be the integral operator on $L^{2}([0,1])$ with kernel $W$ and eigenvalues $\lambda_{1}(T_{W})\geq\lambda_{2}(T_{W})\geq\cdots$. Fix $r$ as in Assumption 1. Then
@@
Lemma 11 (Resolvent correction in the degenerate regime).
@@
Fix the approximate eigenvectors $\mathbf{u}_{k}$ as in Lemma 7, and for each fixed $K$ consider the family $\{\mathbf{u}_{k}\}_{k\leq K,k\neq r}$, which is asymptotically orthonormal and lies in $\mathbf{u}^{\perp}$ up to $o_{p}(1)$ by Lemma 7. By Gram-Schmidt, choose $\mathbf{V}$ so that, for each fixed $K$, its first K1K-1 columns agree with an orthonormalization of $\{\mathbf{u}_{k}\}_{k\leq K,k\neq r}$, noting that the quadratic form
@@
depends only on the subspace $\mathbf{u}^{\perp}$ and $A$, not on the specific orthonormal basis $\mathbf{V}$, so this choice is without loss of generality. By Lemma 10 and the simple eigengap at $\lambda_{r}$, the eigenvalues of $(n-1)^{-1}\mathbf{K}_{n}$ in $\mathbf{u}^{\perp}$ concentrate around $\{\lambda_{k}\}_{k\neq r}$, and hence
@@
so that $s_{k,n}=n(1+o_{p}(1))$ by Lemma 15, and Lemma 7 gives $\mathbf{u}_{k}^{\top}\mathbf{u}_{\ell}=\delta_{k\ell}+o_{p}(1)$ and
@@
where we used lemma  7 for $\mathbf{U}^{\top}\mathbf{U}$ and $\mathbf{M}_{K}$. For $k\leq K$, $k\neq r$, the approximate eigenvalue equation yields
@@
Lemma 12 (Second-order Rayleigh-Schrödinger expansion for a simple eigenvalue).
@@
Lemma 13 (Degenerate U-statistic in the degenerate regime).
@@
Lemma 14 (Square-summability of fluctuation coefficients).
Let $T_{W}:L^{2}([0,1])\to L^{2}([0,1])$ be the compact self-adjoint operator associated with the symmetric kernel $W\in L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$, and let $\{(\lambda_{k},\varphi_{k})\}_{k\geq 1}$ be its orthonormal eigenpairs. Fix a simple eigenvalue $\lambda_{r}$ with eigengap $\gamma_{r}=\inf_{k\neq r}|\lambda_{r}-\lambda_{k}|>0$. Then
$\sum_{k\neq r}\left(\frac{\lambda_{r}\lambda_{k}}{\lambda_{r}-\lambda_{k}}\right)^{2}<\infty.$
Proof.
Since $T_{W}$ is Hilbert-Schmidt, we have $\sum_{k=1}^{\infty}\lambda_{k}^{2}=\|W\|_{L^{2}([0,1]^{2})}^{2}<\infty.$ For $k\neq r$, note that $|\lambda_{r}-\lambda_{k}|\geq\gamma_{r}>0$ by the eigengap assumption, so
$\left(\frac{\lambda_{r}\lambda_{k}}{\lambda_{r}-\lambda_{k}}\right)^{2}\leq\frac{\lambda_{r}^{2}}{\gamma_{r}^{2}}\,\lambda_{k}^{2}.$
Summing over all $k\neq r$ yields
$\sum_{k\neq r}\left(\frac{\lambda_{r}\lambda_{k}}{\lambda_{r}-\lambda_{k}}\right)^{2}\leq\frac{\lambda_{r}^{2}}{\gamma_{r}^{2}}\sum_{k\neq r}\lambda_{k}^{2}\leq\frac{\lambda_{r}^{2}}{\gamma_{r}^{2}}\|W\|_{L^{2}([0,1]^{2})}^{2}<\infty.$
Hence the coefficients are square-summable, and the random series $\sum_{k\neq r}\tfrac{\lambda_{r}\lambda_{k}}{\lambda_{r}-\lambda_{k}}(Z_{k}^{2}-1)$ converges in $L^{2}$ by the completeness of $\ell^{2}$ and independence of the Gaussian sequence $(Z_{k})$. ∎
Lemma 15.